# LoLA: Low-Rank Linear Attention With Sparse Caching

### v1

@@ Appendix B Related Work
State Space Models (SSMs) have emerged as powerful architectures for efficient long-range sequence modeling, offering constant memory complexity irrespective of context length. Pioneered by methods like S4 [22], recent developments include various efficient architectures such as RetNet [44] and Mamba [21]. Concurrently, original linear attention methods have explored efficient approximations of the softmax kernel [27, 11, 37, 36, 57]. Research between SSMs and linear attention has recently converged. Modern SSMs, such as DeltaNet [41], Mamba2 [14] and Gated DeltaNet [51], can be interpreted as linear attention models equipped with additional gating or delta-update mechanisms. Models like DeltaProduct [42] further generalize these linear updates through higher-rank modifications, while Titans [3] and TTT [45] extend the capacity for associative recall using richer hidden-state representations.
Test-time Regression [50] offers a unifying perspective for SSMs and linear attention. These sequence models perform online regression to fit hidden states to past context. Each state update can be interpreted as a gradient step in online SGD. This lens clarifies the roles of different mechanisms within these models. For example, forget gates in Mamba2 and Gated DeltaNet play an analogous role to weight decay. Similarly, momentum-based updates can be oversed in Titans [3].
@@
The cost of pretraining LLMs is the primary obstacle in finding the successor of the transformer. To address this issue, recent approaches employ knowledge distillation, transferring the capabilities of pretrained transformers into subquadratic architectures [6, 56, 30, 5, 19]. This significantly reduces training costs by recycling large pretrained models.
MOHAWK [6, 5] demonstrated successful distillation of pretrained transformers into Mamba, maintaining competitive performance. Similarly, “Mamba in the Llama” [49] interleaves transformer and SSM blocks to retain transformer-level performance with significantly reduced inference costs. Though, this approach maintains an unbounded memory footprint due to the residual quadratic attention.
@@
Sparse attention methods present another orthogonal approach to reducing Transformer complexity by limiting the set of attended tokens [31]. Methods such as Longformer [4] and BigBird [54] adopt fixed sparse patterns that incorporate sliding windows and selective global attention, efficiently capturing both local and sparse global contexts. Recent dynamic sparsification approaches, including Loki [43] and Native Sparse Attention (NSA)[53], employ data-dependent strategies, selectively attending to the most relevant tokens based on learned or projected keys. Native Sparse Attention, specifically, combines sparse attention with latent attention mechanisms[29], effectively approximating attention via low-rank and sparse structures.
We believe sparse attention can be complementary to linear attention. With LoLA, we encourage the use of hybrid attention techniques within the same attention head. This allows important tokens to leverage more computation when needed. This work contrasts the use of interleaving soft attention blocks with linear attention blocks [39, 18] which allocates the compute costs equally between all tokens.
Appendix C Linear Attention is a Bad Low-Rank Approximation
@@
which has a maximum rank of $D$. Ideally, $\widehat{\mathbf{G}}$ would closely approximate $\mathbf{G}$, minimizing the squared Frobenius norm error. However, linear attention’s approximation error is fundamentally lower-bounded by the truncated singular value decomposition (SVD) of $\mathbf{G}$ [16]. Specifically, for the SVD decomposition $\mathbf{G}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$ with singular values $\sigma_{i}$, we have:
@@
Figure 4: Visualization of singular values of the Gram matrix and minimum squared Frobenius norm error for linear attention (17). We vary the number of i.i.d. vectors $n$ used to construct the Gram matrix, but maintain the same input dimension $d$.
Figure 5: Visualization of singular values of the Gram matrix and minimum squared Frobenius norm error for linear attention (17). We vary the input dimension $d$ between columns in the plot.
Figures 4 and 5 show that applying an exponential operation to the query-key products significantly increases the rank and complexity of the resulting Gram matrix. The singular values and approximation error increase with the number of unique input vectors $n$ (see Figure 4) and the input dimension $d$ (see Figure 5) . Practically, in transformer architectures, head dimensions are typically modest ($d=64$ for Llama-3.2 1B and $d=128$ for Llama-3.1 8B). Additionally, linear attention approaches typically select feature dimensions $D$ around $2d$ [57, 56] which can be problematic without additional sparse attention or gating mechanisms.
@@
Appendix D Scoring Ablation Extension
In section 4.3, we ablated different scoring approaches for the sparse cache. Here, we describe exactly how each alternative score is computed. As a reminder, LoLA is motivated by the occurence of memory collisions in the hidden state. LoLA explicitly attempts to maintain self-recall for stored key-value pairs. An alternative perspective could be aiming for the best softmax approximation. Similar to previous kernel work [11], this aims to minimize the attention weight error
@@ Appendix D Scoring Ablation Extension
Appendix E Extended Memory Collision Visualization
Figure 6: Visualizing the relative self-recall error for stored KV pairs.
In Section 4.3, we visualized how memory collisions can occur in practice. This was computed by measuring the self-recall error for all stored KV pairs. We found that the first few stored KV pairs do not achieve a large error when added to the hidden state. However, these quickly become corrupted after hidden state updates. This phenomenon was difficult to see in Figure 3, so we provide an additional visualization with Figure 6. Specifically, we measure the self-recall error of each KV pair, relative to the error of when that pair was added. For row $i$ column $j$ in the plot, the relative error is computed as
@@ Appendix E Extended Memory Collision Visualization
Here, we see the first few KV pairs have high relative errors for pure linear attention. These pairs observe small self-recall errors at early time steps, but achieve much higher self-recall errors later. On the other hand, sparse caching actively mitigates self-recall errors, improving associative recall for stored KV pairs.
Appendix F Extended Long Context Tasks
To further extend our results on long-context recall tasks, we report scores for the rest of the RULER benchmark [25] in table 7. These represent harder, multi-step tasks: Needle-in-a-Haystack-style Multi-Key Retrieval (MK1, MK2, MK3), NIAH-style Multi-Query (MQ), NIAH Multi-Value (MV), Variable Tracking (VT), Commonword Extraction (CWE), Frequent Word Extraction (FWE), HotpotQA, and SquadQA. Due to the difficulty of these tasks, we benchmark LoLA on 2048 sequence lengths when applicable. Otherwise, we use the default 4096.
Table 7: Extended results on the RULER benchmark. LoLA $4+$ uses chunkwise inference with $C=\lambda=512$, similar to Table 2.
Task MK1 MK2 MK3 MQ MV VT CWE FWE HotpotQA SquadQA
Seq. Len 2K 2K 2K 2K 2K 4K 4K 4K 4K 4K
LoLA $4+$ 70.0 20.2 0 62.75 54.15 8.3 1.9 12.93 25.2 51.7
Appendix G Algorithm Psuedo-code
We provide PyTorch-like pseudo-code for the cache during decoding or generation. In this example, we update the cache every iteration (rather than chunkwise inference) for simplicity. Lastly, this psuedo-code does not contain any optimization tricks for ease of understanding.

### v2

@@ Appendix D Related Work
State Space Models (SSMs) have emerged as powerful architectures for efficient long-range sequence modeling, offering constant memory complexity irrespective of context length. Pioneered by methods like S4 (Gu et al., 2021), recent developments include various efficient architectures such as RetNet (Sun et al., 2023) and Mamba (Gu & Dao, 2024). Concurrently, original linear attention methods have explored efficient approximations of the softmax kernel (Katharopoulos et al., 2020; Choromanski et al., 2020; Qin et al., ; Peng et al., 2020; Zhang et al., 2024). Research between SSMs and linear attention has recently converged. Modern SSMs, such as DeltaNet (Schlag et al., 2021), Mamba2 (Dao & Gu, 2024) and Gated DeltaNet (Yang et al., 2024a), can be interpreted as linear attention models equipped with additional gating or delta-update mechanisms. Models like DeltaProduct (Siems et al., 2025) further generalize these linear updates through higher-rank modifications, while Titans (Behrouz et al., 2024) and TTT (Sun et al., 2024) extend the capacity for associative recall using richer hidden-state representations.
Test-time Regression (Wang et al., 2025) offers a unifying perspective for SSMs and linear attention. These sequence models perform online regression to fit hidden states to past context. Each state update can be interpreted as a gradient step in online SGD. This lens clarifies the roles of different mechanisms within these models. For example, forget gates in Mamba2 and Gated DeltaNet play an analogous role to weight decay. Similarly, momentum-based updates can be oversed in Titans (Behrouz et al., 2024).
@@
The cost of pretraining LLMs is the primary obstacle in finding the successor of the transformer. To address this issue, recent approaches employ knowledge distillation, transferring the capabilities of pretrained transformers into subquadratic architectures (Bick et al., 2024; Zhang et al., 2025a; Mercat et al., ; Bick et al., 2025; Goldstein et al., 2025). This significantly reduces training costs by recycling large pretrained models.
MOHAWK (Bick et al., 2024; 2025) demonstrated successful distillation of pretrained transformers into Mamba, maintaining competitive performance. Similarly, “Mamba in the Llama” (Wang et al., 2024) interleaves transformer and SSM blocks to retain transformer-level performance with significantly reduced inference costs. Though, this approach maintains an unbounded memory footprint due to the residual quadratic attention.
@@
Sparse attention methods present another orthogonal approach to reducing Transformer complexity by limiting the set of attended tokens (Nawrot et al., 2025). Methods such as Longformer (Beltagy et al., 2020) and BigBird (Zaheer et al., 2020) adopt fixed sparse patterns that incorporate sliding windows and selective global attention, efficiently capturing both local and sparse global contexts. Recent dynamic sparsification approaches, including Loki (Singhania et al., 2024) and Native Sparse Attention (NSA)(Yuan et al., 2025), employ data-dependent strategies, selectively attending to the most relevant tokens based on learned or projected keys. Native Sparse Attention, specifically, combines sparse attention with latent attention mechanisms(Liu et al., 2024), effectively approximating attention via low-rank and sparse structures.
We believe sparse attention can be complementary to linear attention. With LoLA, we encourage the use of hybrid attention techniques within the same attention head. This allows important tokens to leverage more computation when needed. This work contrasts the use of interleaving soft attention blocks with linear attention blocks (Ren et al., 2025; Glorioso et al., 2024) which allocates the compute costs equally between all tokens.
Appendix E Linear Attention is a Bad Low-Rank Approximation
@@
which has a maximum rank of $D$. Ideally, $\widehat{\mathbf{G}}$ would closely approximate $\mathbf{G}$, minimizing the squared Frobenius norm error. However, linear attention’s approximation error is fundamentally lower-bounded by the truncated singular value decomposition (SVD) of $\mathbf{G}$ (Eckart & Young, 1936). Specifically, for the SVD decomposition $\mathbf{G}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$ with singular values $\sigma_{i}$, we have:
@@
Figure 8: Visualization of singular values of the Gram matrix and minimum squared Frobenius norm error for linear attention as in equation 18. We vary the number of i.i.d. vectors $n$ used to construct the Gram matrix, but maintain the same input dimension $d$.
Figure 9: Visualization of singular values of the Gram matrix and minimum squared Frobenius norm error for linear attention as in equation 18. We vary the input dimension $d$ between columns in the plot.
Figures 8 and 9 show that applying an exponential operation to the query-key products significantly increases the rank and complexity of the resulting Gram matrix. The singular values and approximation error increase with the number of unique input vectors $n$ (see Figure 8) and the input dimension $d$ (see Figure 9) . Practically, in transformer architectures, head dimensions are typically modest ($d=64$ for Llama-3.2 1B and $d=128$ for Llama-3.1 8B). Additionally, linear attention approaches typically select feature dimensions $D$ around $2d$ (Zhang et al., 2024; 2025a) which can be problematic without additional sparse attention or gating mechanisms.
@@
Appendix F Scoring Ablation Extension
In section 4.3, we ablated different scoring approaches for the sparse cache. Here, we describe exactly how each alternative score is computed. As a reminder, LoLA is motivated by the occurence of memory collisions in the hidden state. LoLA explicitly attempts to maintain self-recall for stored key-value pairs. An alternative perspective could be aiming for the best softmax approximation. Similar to previous kernel work (Choromanski et al., 2020), this aims to minimize the attention weight error
@@ Appendix F Scoring Ablation Extension
From traditional sparse attention literature (Zhang et al., 2023; Dong et al., 2024), keys that are highly attended to may be important. We compute this as
$\text{Score}(\bm{k}_{i})=\sum_{t=i}^{i+\eta-1}\exp(\bm{q}_{t}^{\top}\bm{k}_{i}).$ (21)
@@
Appendix G Extended Memory Collision Visualization
Figure 10: Visualizing the relative SRE for stored KV pairs.
In Section 4.3, we visualized how memory collisions can occur in practice. This was computed by measuring the SRE for all stored KV pairs. We found that the first few stored KV pairs do not achieve a large error when added to the hidden state. However, these quickly become corrupted after hidden state updates. This phenomenon was difficult to see in Figure 3, so we provide an additional visualization with Figure 10. Specifically, we measure the SRE of each KV pair, relative to the error of when that pair was added. For row $i$ column $j$ in the plot, the relative error is computed as
@@ Appendix G Extended Memory Collision Visualization
Here, we see the first few KV pairs have high relative errors for pure linear attention. These pairs observe small SREs at early time steps, but achieve much higher SREs later. On the other hand, sparse caching actively mitigates SREs, improving associative recall for stored KV pairs.
Appendix H Algorithm Psuedo-code
We provide PyTorch-like pseudo-code for the cache during decoding or generation. In this example, we update the cache every iteration (rather than chunkwise inference) for simplicity. Lastly, this pseudo-code does not contain any optimization tricks for ease of understanding.